OpenStudy (anonymous):

an equation for a line is 3x - 4y = 8 what is the slope of a line perpendicular to the given line

Nnesha (nnesha):

first solve this one for y in other words set this to slope intercept form

OpenStudy (anonymous):

can you help me with the answer

Nnesha (nnesha):

first solve then we get answer sound good to me right ????

OpenStudy (anonymous):

but i dont know how

Nnesha (nnesha):

ok no problem :) that's why i'm here to help :)

Nnesha (nnesha):

ok to solve for y we have to move 3x first 3x is adding on left side what is opposite of add ??

OpenStudy (anonymous):

the answer choices are A. 6x - 5y = 17 . then B . 6x + 5y = 7 then C . 5x - 6y = - 45 and D is 5x + 6y = 4

OpenStudy (anonymous):

the opposite of add is subtract

Nnesha (nnesha):

yes so subtract 3x both side  now let me know what you get after subtrracting